# Active Learning of A Crystal Plasticity Flow Rule From Discrete Dislocation Dynamics Simulations

### v1

@@ 1 Introduction
Dislocation-mediated plasticity is the primary mechanism for inelastic deformation in crystalline materials. Computational methods of predicting deformation of single- and poly-crystal materials are of great importance in a multitude of structural applications. A commonly used computational method is Crystal Plasticity (CP) which predicts continuum-scale deformation using constitutive representations of the various deformation phenomena occurring at lower scales using a so-called flow rule. The flow rule predicts single crystal strain rates along crystallographic slip systems in response to an applied load, encapsulating complex motions and interactions of dislocations as a function of dislocation density and stress, see, e.g., Ref. [1].
@@
To date, flow rules in crystal plasticity have primarily, if not entirely, taken on forms based upon phenomenology or constitutive laws. References [2] and [3] illustrate and address the challenges in designing a flow rule which accurately represents the deformation dynamics of BCC structured materials using phenomenologically based formulae. Some of these challenges may be addressed by inspecting simulations of BCC deformation dynamics at the atomic scale, and developing customized flow rule formulae to that data, as in [4] and [5]. The drawback to this approach is that atomic simulations of sufficient size are so costly that opportunities to produce a wide enough variety of them to extract parameters of a flow rule are uncommon.
@@
To avoid the high cost of atomic simulations, Discrete Dislocation Dynamics (DDD) simulations aim to predict crystalline deformation directly at the meso-scale by representing dislocations as simple line defects whose interactions are mediated by linear elasticity. By enabling a significantly larger set of training data at sufficient size, DDD simulations can provide the physical insight required to propose corrections to legacy phenomenological relationships for plastic flow and dislocation multiplication, as well as to develop and fit new constitutive flow rules [6].
@@
Prior work along these lines used data generated from simulations or experiments whose parameters were chosen heuristically, producing an expert’s best guess at an adequate representation of the parameter space of applied mechanical loads and dislocation configurations [6, 7, 2, 4, 8, 5]. Generating the corresponding data required human expertise and labor. To expedite and reduce the cost of generating the data required for fitting a model, it is highly desirable to automate the simulations through methods which optimize the parameter space exploration aided by an uncertainty metric directly embedded in the model. In this work, we demonstrate a semi-autonomous computational pipeline which iteratively instantiates a sequence of DDD simulations, efficiently navigating the parameter space of a flow rule to build a data set sufficient to fit a crystal plasticity flow rule.
@@
The manuscript is organized as follows. Section 2 discusses in detail our choices of: mesoscale model to be fitted and parameterization of its domain (2.1), the method by which this model is used to identify simulation specifications to be instantiated (2.2), the constraints used to ensure physical and practical plausibility (2.3), the implementation of DDD used (2.4), processing and preparation of the simulation measurements to be presented to the model (2.5), and the overall workflow and automation methods we implemented (2.6). In Section 3, we present examples of pertinent measurements from DDD (3.1), and an assessment of the quality of the model as it evolved through the sequence of simulations which it guided (3.2). Finally, we briefly summarize our achievements and conclude in Section 4. Several appendices are provided as supplementary information.
@@ 2.1 Model and optimization
Sequential exploration of a complex parameter space benefits from insight from the model to guide its own improvement using a local measure of uncertainty. The high cost of simulations and thus likely restrictions on data set size must also be considered, as many candidate machine learning methods require a prohibitively large amount of data. In this context we prioritize optimizing the efficiency with which we navigate the high dimensional parameter space to gather data which is most effective at reducing predictive errors and at naturally encapsulating a wide variety of the underlying physical phenomena.
@@
Gaussian processes (GP) have been broadly utilized in situations like these. A GP is a collection of random variables, any finite subset of which follows a joint multivariate Gaussian distribution. It defines a prior over functions, which can be updated to a posterior given observed data, making it a non-parametric Bayesian method. A GP is fully specified by a mean function and a covariance (kernel) function, allowing for flexible non-parametric modeling of complex functions. It also naturally provides uncertainty quantification through estimates of the variance of the posterior distribution. An effective active learning scheme is then achieved by instantiating simulations at points for which the model’s posterior distribution has a high variance, as chosen by a so-called acquisition function, with the expectation that obtaining additional data at those points will reduce the said variance and thus the local model uncertainty. Such a method of active learning that implements GP regression (GPR) in this way is commonly known as Bayesian optimization (BO) [9].
@@
Formulation of the GPR model and its domain (i.e., input) variables should be sufficiently complex to capture and describe the target physical phenomenon, but should otherwise be kept as compact as possible, as it has been well established that the efficiency of GPR tend to degrade in dimensions much higher than about 10 [10].
@@
In our context, dislocation densities of the twenty four 111{110} slip systems are the obvious candidates for dimensional reduction for several reasons. While coupling between densities of separate slip systems is known to have a critical influence on slip system strain rates[11, 12], and thus should not be omitted entirely from a flow rule model domain, a full specification of the densities on every slip system is not necessary to characterize these variations. Conventional methods of reducing the number of slip system densities involve phenomenologically-derived coupling constants which are used to calculate lower dimensional representations of the complete density specification in the frame of a given slip system [8, 6].
@@
With these issues in mind, we’ve chosen to restrict the model’s inputs to only two dislocation density parameters, being that of a particular 111{110} slip system and the total dislocation density over all slip systems . This corresponds to the simplest form of flow-rule used in Crystal Plasticity models, encoding Taylor hardening and Kocks-Mecking formalism for evolution of dislocation density. This choice preserves the ability of the acquisition function to suggest configurations which may have strong coupling between slip system dislocation densities, in which is comparably greater than , or configurations dominated by a single slip system where is approximately , while reducing the total volume of the parameter space. It should be noted, however, that reducing completeness of the specification by combining dislocation densities into a single parameter potentially induces noise as there will multiple ways to instantiate the same total dislocation density, each of which could yield a different strain-rate.
@@
The remaining simulation parameters thought to determine strain rate are those describing the mechanical load which are summarized by the six unique applied stress tensor components, , , , , , , defined with respect to a local coordinate system aligned with the slip plane and the slip direction to comply with equivalence among 111{110} slip systems. Retaining all six of these dimensions as free parameters, which may take on either positive or negative values, allows the model to capture the non-Schmid effects observed in experiments[13, 14] and predicted in molecular dynamics simulations of BCC materials [5, 15, 16]. Our final GPR model is hence formulated using an eight-dimensional slip-system specific domain parameterization: where the density is specific to a particular slip system, and stress tensor is a rotation of the global stress into a coordinate system corresponding to that particular slip system in which the slip system’s normalized Burgers vector and slip plane normal vector serve as the first and third vectors of an orthonormal basis, respectively. The slip systems are enumerated in Table 1 and illustrated in Figure 9.
@@
The GP is implemented using the SingleTaskGP model from BoTorch using a Matèrn covariance kernel as implemented by GPyTorch with homoscedastic noise and fixed smoothness parameter =2.5, along with a Gaussian likelihood with gamma prior. Fitting of the model to training data is performed after first initializing length-scale hyper-parameters to values representative of the anticipated sensitivities of the objective to changes in corresponding model domain parameters. For example, strain rate is anticipated to be highly sensitive to resolved shear stress as represented by the stress component in the model’s rotated frame and to the dislocation density of the given slip system, and thus ought to have shorter length scale hyper-parameters along those dimensions. During model fitting, these initial values are optimized along with a scalar noise hyper-parameter which embodies a prediction of variance or noise in the outcome and is reported as a predictive likelihood variance or standard deviation in what follows. Allowing hyper-parameters to be optimized during the fitting procedure in this way is sometimes referred to as the empirical Bayes approach [17, 18]. Having chosen to use homoscedastic noise, the noise hyper-parameter, length-scale hyper-parameters, and likelihood variance will be scalars which do not vary within the input domain.
@@ 2.2 Acquisition function
The optimal microstructural and loading parameters used to instantiate batches of simulations are sequentially identified by a custom-designed acquisition function. This acquisition function is designed to favor parameters where explicit simulation results are anticipated to maximally reduce the posterior variance. A second requirement is the ability to identify multiple simulations (i.e., a so-called batch in the BO parlance) that can be executed concurrently while avoiding excessive inter-task redundancy. This batched mode of operation enable the reduction in the overall time needed to obtain the required simulation data through the use of parallel computing resources. To guide our simulations with these issues in mind, we’ve implemented an acquisition function which adapts the idea of reconditioning a temporary model on its own predictions, creating a so-called fantasy model [19, 20]. In this approach, the model used for BO is iteratively retrained using virtual observations that are represented as the posterior mean prediction of the current model. This has the effect of decreasing the estimated posterior variance around the virtual observation, discouraging the acquisition of new measurements in the neighborhood of the virtual measurement. The batch (with simulations) is then constructed by sequentially applying the acquisition function on the reconditioned fantasy model to produce a batch of experiments that are globally maximally informative and non-redundant.
A non-trivial challenge in our context is to design an acquisition function which accommodates the difference in dimensionality between the eight dimensional GP model domain and the thirty dimensional DDD simulation domain consisting of microstructural and loading parameters. Due to the equivalence of the twenty four slip systems in BCC crystals, each point in the DDD simulation domain uniquely determines a collection of twenty four points in the GP model domain, one for each slip system. This equivalence places constraints between the members of each subset of model domain points that correspond to a single simulation domain point, specifically that they must all have stress tensors which are equivalent through a transformation from the global simulation frame to their individual slip system frame: , where is the rotation matrix from a global Cartesian coordinate system to the coordinate system having basis . Secondly, the sum of all slip system densities within the simulation domain must be equal to the total density parameter of the model domain: .
In addition to the above constraints, required by the dimensionality mismatch between model and simulation, a mechanism to prevent degeneracy between model domain points is also desired. For this we’ve implemented a method of sequentially identifying points in the model domain under constraints which evolve with each proposed candidate. This is achieved by reconditioning the model on every 8-D candidate as soon as they are proposed, using the model’s posterior mean as if it were an accurate prediction. This reconditioning temporarily reduces the model variance in the neighborhood of those proposed points, thus reducing the propensity for the acquisition function to repeatedly propose similar points within a single batch [19, 20]. By preserving the fantasized model throughout the loop which identifies a single batch, similarity is dissuaded between all of the corresponding 8-D model domain points of the entire batch, considering all slip systems within the same configuration as well as slip systems present in different configurations that are part of the same batch. The flow of the acquisition function is outlined in Algorithm 1.
@@
Unlike conventional Bayesian optimization in which the goal is to identify the maximum value of an expensive function, our goal is to optimize the fitting of a Gaussian process model over a prescribed domain. Conventional acquisition functions like the well known expected improvement (EI) or knowledge gradient (KG) are thus not easily adaptable. Instead, by assuming that the variance at each model domain point is a measure of model uncertainty, we seek new training data at points where the GP model predicts highest variance, which has been argued to be equivalent to reducing the Shannon entropy of the random variable which the Bayesian model represents and has been previously applied to optimal design of computer experiments [21, 22].
@@
Due to the dimensionality and potential roughness of the search landscape, a multi-start optimization scheme from the BoTorch software package [23] is used in which collections of random starting points are produced via Monte Carlo Sampling (MCS). Local maxima of model posterior variance near each starting point are then found via the sequential least squares programming (SLSQP) method, as implemented in the SciPy software package [24]. After repeating the multistart optimization a specified number of times and selecting the point for which the model posterior has maximum variance within the results, the point representing the first slip system in the model domain is obtained. The fantasized model is then reconditioned on this point using the value of its own posterior mean. During this routine of proposing and optimizing candidates in the model domain, constraints discussed in the next section are enforced to ensure that the candidates are consistent with physical and practical limitations. All but the final of the remaining slip system density specifications are then determined sequentially using the same multi-start optimization method but under repeatedly updated constraints and reconditioning of the fantasized model. These constraints restrict model domain candidates to stress tensors which are transformed by the aforementioned rotations and values of density that must lie between a minimum density and an appropriate fraction of the remaining available density, for the slip system where , to allow for each remaining slip system to be assigned a density of at least . The final slip system is eventually assigned all of the remaining available density, resulting in a collection of densities whose total was determined by the first step. In this way all thirty parameters in the domain of a single simulation target are chosen: . The temporary reconditioned fantasy model is preserved until all targets in a batch of simulations are specified, to encourage variety within each batch.

### v2

@@ 1 Introduction
Dislocation-mediated plasticity is the primary mechanism for inelastic deformation in crystalline materials. Computational methods of predicting deformation of single- and poly-crystal materials are of great importance in a multitude of structural applications. A commonly used computational method is Crystal Plasticity (CP) which predicts continuum-scale deformation using constitutive representations of the various deformation phenomena occurring at lower scales using a so-called flow rule. The flow rule predicts single crystal strain rates along crystallographic slip systems in response to an applied load, encapsulating complex motions and interactions of dislocations as a function of dislocation density and stress, see, e.g., Ref. [ROTERS20101152].
@@
To date, flow rules in crystal plasticity have primarily, if not entirely, taken on forms based upon phenomenology or constitutive laws. References [MA200791] and [CHO2018138] illustrate and address the challenges in designing a flow rule which accurately represents the deformation dynamics of BCC structured materials using phenomenologically based formulae. Some of these challenges may be addressed by inspecting simulations of BCC deformation dynamics at the atomic scale, and developing customized flow rule formulae to that data, as in [LIM2015100] and [BERTIN2023119336]. The drawback to this approach is that atomic simulations of sufficient size are so costly that opportunities to produce a wide enough variety of them to extract parameters of a flow rule are uncommon.
@@
To avoid the high cost of atomic simulations, Discrete Dislocation Dynamics (DDD) simulations aim to predict crystalline deformation directly at the meso-scale by representing dislocations as simple line defects whose interactions are mediated by linear elasticity. By enabling a significantly larger set of training data at sufficient size, DDD simulations can provide the physical insight required to propose corrections to legacy phenomenological relationships for plastic flow and dislocation multiplication, as well as to develop and fit new constitutive flow rules [AKHONDZADEH2020104152].
@@
Prior work along these lines used data generated from simulations or experiments whose parameters were chosen heuristically, producing an expert’s best guess at an adequate representation of the parameter space of applied mechanical loads and dislocation configurations [AKHONDZADEH2020104152, MA20043603, MA200791, LIM2015100, LEE1999625, BERTIN2023119336].Generating the corresponding data required human expertise and labor. To expedite and reduce the cost of generating the data required for fitting a model, it is highly desirable to automate the simulations through methods which optimize the parameter space exploration aided by an uncertainty metric directly embedded in the model. In this work, we demonstrate a semi-autonomous computational pipeline which iteratively instantiates a sequence of DDD simulations, efficiently navigating the parameter space of a flow rule to build a data set sufficient to fit a crystal plasticity flow rule.
@@
The manuscript is organized as follows. Section 2 discusses in detail our choices of: mesoscale model to be fitted (2.1), the parameterization of the model domain (2.2), implementation of the chosen model form (2.3), the method by which this model is used to identify simulation specifications to be instantiated (2.4), the constraints used to ensure physical and practical plausibility (2.5), the implementation of DDD used (2.6), processing and preparation of the simulation measurements to be presented to the model (2.7), and the overall workflow and automation methods we implemented (2.8). In Section 3, we present examples of pertinent measurements from DDD (3.1), and an assessment of the quality of the model as it evolved through the sequence of simulations which it guided (3.2). Finally, we briefly summarize our achievements and conclude in Section 4. Several appendices are provided as supplementary information.
@@ 2.1 Model selection
Sequential exploration of a complex parameter space may benefit from insight from the model to guide its own improvement using a local measure of uncertainty. The high cost of simulations and thus likely restrictions on data set size must also be considered, as many candidate machine learning methods require a prohibitively large amount of data. When choosing a model, its parameterization, and corresponding transformations to and from DDD simulation data in this context we prioritize optimizing the efficiency with which we navigate the high dimensional parameter space to gather data which is most effective at reducing predictive errors and at naturally encapsulating the basic underlying physical phenomena. The data generated may be used in future work to fit models whose parameterization and fidelity may be tuned to optimize behavior in a CP context.
@@
Gaussian processes (GP) have been broadly utilized in situations like these. Abstractly, a GP is a collection of random variables, any finite subset of which follows a joint multivariate Gaussian distribution. As a consequence, an instance of a GP is fully specified by a mean function and a covariance (kernel) function which may be fitted to data by optimizing their hyperparameters, allowing for flexible modeling of complex functions. Additionally, GP models are well suited to be sequentially updated as new information becomes available through the application of Bayes’ Rule since the initial state, called the prior, the prediction of the model, called the likelihood, and the resulting updated distribution, called the posterior, have easily obtainable formulae. The variance of the posterior is then a natural quantification of the uncertainty of the model. An effective active learning scheme is then achieved by instantiating simulations at points for which the model’s posterior distribution has a high variance, as chosen by a so-called acquisition function, with the expectation that obtaining additional data at those points will reduce the said variance and thus the local model uncertainty. Such a method of active learning that implements GP regression (GPR) in this way is commonly known as Bayesian optimization (BO) [Frazier2018].
2.2 Model parameterization
@@
Formulation of the GPR model and its domain (i.e., input) variables should be sufficiently complex to capture and describe the target physical phenomenon, but should otherwise be kept as compact as possible, as it has been well established that the efficiency of GPR tend to degrade in dimensions much higher than about 10 [10.1145/3545611].
@@
In our context, dislocation densities representing the total lengths of dislocations per unit volume of the twenty four 111{110} slip systems are the obvious candidates for dimensional reduction for several reasons. While coupling between densities of separate slip systems is known to have a critical influence on slip system strain rates[BULATOV20255611, akhondzadeh2021slip], and thus should not be omitted entirely from a flow rule model domain, a full specification of the densities on every slip system is not necessary to characterize these variations. Conventional methods of reducing the number of slip system densities involve phenomenologically-derived coupling constants which are used to calculate lower dimensional representations of the complete density specification in the frame of a given slip system[LEE1999625, AKHONDZADEH2020104152].
@@
With these issues in mind, we’ve chosen to restrict the model’s inputs to only two dislocation density parameters, being that of a particular 111{110} slip system and the total dislocation density over all slip systems . This corresponds to the simplest form of flow-rule used in Crystal Plasticity models, encoding Taylor hardening [taylor1934mechanism] Kocks-Mecking formalism [kocks2003physics] for evolution of dislocation density. This choice preserves the ability of the acquisition function to suggest configurations which may have strong coupling between slip system dislocation densities, in which is comparably greater than , or configurations dominated by a single slip system where is approximately , while reducing the total volume of the parameter space. It should be noted, however, that reducing completeness of the specification by combining dislocation densities into a single parameter potentially induces noise as there will multiple ways to instantiate the same total dislocation density, each of which could yield a different strain-rate.
@@
The remaining simulation parameters thought to determine strain rate are those describing the mechanical load which are summarized by the six unique applied stress tensor components, , , , , , , defined with respect to a local coordinate system aligned with the slip plane and the slip direction to comply with equivalence among 111{110} slip systems. Retaining all six of these dimensions as free parameters, which may take on either positive or negative values, allows the model to capture the non-Schmid effects observed in experiments[christian1970dislocations, christian1983some] and predicted in molecular dynamics simulations of BCC materials [BERTIN2023119336, Kraych2019, Dezerald_2016]. Our final GPR model is hence formulated using an eight-dimensional slip-system specific domain parameterization: where the density is specific to a particular slip system, and stress tensor is a rotation of the global stress into a coordinate system corresponding to that particular slip system in which the slip system’s normalized Burgers vector and slip plane normal vector serve as the first and third vectors of an orthonormal basis, respectively. The slip systems are enumerated in Table 1 and illustrated in Figure 9.
2.3 Details of the GP implementation
@@
The GP is implemented using the SingleTaskGP model from BoTorch using a Matèrn covariance kernel as implemented by GPyTorch with homoscedastic noise and fixed smoothness parameter =2.5, along with a Gaussian likelihood with gamma prior. Fitting of the model to training data is performed after first initializing length-scale hyper-parameters to values representative of the anticipated sensitivities of the objective to changes in corresponding model domain parameters. For example, strain rate is anticipated to be highly sensitive to resolved shear stress as represented by the stress component in the model’s rotated frame and to the dislocation density of the given slip system, and thus ought to have shorter length scale hyper-parameters along those dimensions. During model fitting, these initial values are optimized along with a scalar noise hyper-parameter which embodies a prediction of variance or noise in the outcome and is reported as a predictive likelihood variance or standard deviation in what follows. Allowing hyper-parameters to be optimized during the fitting procedure in this way is sometimes referred to as the empirical Bayes approach [dunlop2020hyperparameter, doi:10.1137/19M1284816]. Having chosen to use homoscedastic noise, the noise hyper-parameter, length-scale hyper-parameters, and likelihood variance will be scalars which do not vary within the input domain.
@@ 2.4 Acquisition function
Determining optimal microstructural and loading parameters used to instantiate simulations is the task of a custom-designed acquisition function. This acquisition function is designed to sequentially identify simulation parameters where explicit simulation results are anticipated to maximally reduce the model uncertainty, which is equated with the GP model’s posterior variance as described in Section 2.1. A second requirement is the ability to identify multiple simulations (i.e., a so-called batch in the BO parlance) that can be executed concurrently while avoiding excessive inter-task redundancy. This batched mode of operation enables reduction in the overall time needed to obtain the required simulation data through the use of parallel computing resources.
@@
Unlike conventional Bayesian optimization in which the goal is to identify the maximum value of an expensive function, our goal is to optimize the fitting of a Gaussian process model over a prescribed domain by reducing its uncertainty. Conventional acquisition functions like the well known expected improvement (EI) or knowledge gradient (KG) are thus not easily adaptable. Instead, to reduce the model’s total uncertainty we seek new training data at points where the GP model has highest posterior variance. Updating the model with new information at these points has been argued to be equivalent to reducing the Shannon entropy of the random variable which the model represents and has been previously applied in other works of optimal design of computer experiments [currin1991bayesian, shewry1987maximum].
A non-trivial challenge that arises in our context is to design an acquisition function which accommodates the difference in dimensionality between the eight dimensional GP model domain and the thirty dimensional DDD simulation domain consisting of microstructural and loading parameters. Due to the equivalence of the twenty four slip systems in BCC crystals, each point in the DDD simulation domain uniquely determines a collection of twenty four points in the GP model domain, one for each slip system. This equivalence places constraints between the members of each subset of model domain points that correspond to a single simulation domain point, specifically that they must all have stress tensors which are equivalent through a transformation from the global simulation frame to their individual slip system frame: , where is the rotation matrix from a global Cartesian coordinate system to the coordinate system having basis . Secondly, the sum of all slip system densities within the simulation domain must be equal to the total density parameter of the model domain: .
In addition to the above constraints, required by the dimensionality mismatch between model and simulation, a mechanism to prevent degeneracy between model domain points is also desired. For this, we’ve implemented a method of sequentially identifying points in the model domain under constraints which evolve with each proposed candidate. This is achieved by reconditioning the model on every 8-D candidate as soon as they are proposed, using the model’s posterior mean as if it were an accurate prediction. This reconditioning temporarily reduces the model posterior variance in the neighborhood of those proposed points, thus reducing the propensity for the acquisition function to repeatedly propose similar points within a single batch [NEURIPS2018_498f2c21, lee2020costaware]. By preserving the fantasized model throughout the loop which identifies a single batch, similarity is dissuaded between all of the corresponding 8-D model domain points of the entire batch, considering all slip systems within the same configuration as well as slip systems present in different configurations that are part of the same batch to produce a batch of experiments that are globally maximally informative and non-redundant. The flow of the acquisition function is outlined in Algorithm 1.
@@
Due to the dimensionality and potential roughness of the search landscape, a multi-start optimization scheme from the BoTorch software package [DBLP:journals/corr/abs-1910-06403] is used in which collections of random starting points are produced via Monte Carlo Sampling (MCS). Local maxima of model posterior variance near each starting point are then found via the sequential least squares programming (SLSQP) method, as implemented in the SciPy software package [2020SciPy-NMeth]. After repeating the multistart optimization a specified number of times and selecting the point for which the model posterior has maximum variance within the results, the point representing the first slip system in the model domain is obtained. The fantasized model is then reconditioned on this point using the value of its own posterior mean. During this routine of proposing and optimizing candidates in the model domain, constraints discussed in the next section are enforced to ensure that the candidates are consistent with physical and practical limitations. All but the final of the remaining slip system density specifications are then determined sequentially using the same multi-start optimization method but under repeatedly updated constraints and reconditioning of the fantasized model. These constraints restrict model domain candidates to stress tensors which are transformed by the aforementioned rotations and values of density that must lie between a minimum density and an appropriate fraction of the remaining available density, for the slip system where , to allow for each remaining slip system to be assigned a density of at least . The final slip system is eventually assigned all of the remaining available density, resulting in a collection of densities whose total was determined by the first step. In this way all thirty parameters in the domain of a single simulation target are chosen: . The temporary reconditioned fantasy model is preserved until all targets in a batch of simulations are specified, to encourage variety within each batch.